# Gagan Bhatia ξ El Moatez Billah Nagoudi ξ Abdellah El Mekki ω Fakhraddin Alwajih ξ Muhammad Abdul-Mageedξ,ω,λ

### v2

@@ 4 ArabicMTEB Benchmark
Task Datasets Languages Dialects Metric
@@
In this section, we present ArabicMTEB, a comprehensive Arabic-centric text embedding benchmark designed to evaluate text embeddings across a wide range of tasks and scenarios. ArabicMTEB addresses the limitations of existing benchmarks that either exclude Arabic or lack coverage of diverse Arabic language varieties, dialects, and cultural nuances. Our benchmark includes $94$ datasets spanning $8$ distinct tasks, as summarized in Table 3. Further details about the datasets used in  benchmark  can be found in Appendix C. ArabicMTEB was developed to provide comprehensive coverage of Arabic text embedding capabilities, ensuring the inclusion of MSA and other varieties. It offers diverse task types, such as retrieval, classification, and semantic similarity, to evaluate embeddings holistically across different scenarios by incorporating novel domain-specific, dialectal, and country-level culturally aware datasets, ArabicMTEB represents a more applicable and realistic assessment of Arabic text embeddings.
@@ 4.1 Task Categories
ArabicMTEB categorizes evaluation datasets into the following key task categories, with each type providing a unique perspective on the capabilities of text embeddings. The corresponding metadata for each task, covering the considered number of datasets, number of languages, number of dialects, and evaluation metric, is presented in Table 3:
@@
Semantic Textual Similarity (STS). STS measures the correlation between the embeddings of two sentences, assessing their semantic similarity. ArabicMTEB includes five STS datasets like STS17 and STS22 Cer et al. (2017), along with two synthetic datasets generated using GPT-4 (Details of the creation of these datasets can be found in Appendix I).
@@
Clustering. Clustering groups sentences into clusters based on embedding similarity, evaluating unsupervised learning performance. ArabicMTEB includes four clustering datasets, such as Arabic News Articles and stance detection datasets from baly-etal-2018-integrating.
@@ 4.2 Dialectal ArabicMTEB
Dialectal ArabicMTEB is a specialized fork of the original ArabicMTEB, focusing exclusively on Arabic dialectal datasets. This extension addresses the unique challenges posed by the significant variations in Arabic dialects across different regions, which have been underrepresented in NLP research. While research on dialectal Arabic text embedding has been limited, dialectal ArabicMTEB fills this gap by providing a comprehensive collection of $19$ datasets specifically curated to evaluate embeddings’ performance on diverse Arabic dialects. These datasets span multiple tasks, offering a robust framework for assessing model performance across various dialectal contexts: Bitext Mining: Eight datasets covering dialects such as Algerian, Egyptian, Jordanian, Lebanese, Moroccan, Saudi, and Yemeni Nagoudi et al. (2022); Bouamor et al. (2014). Retrieval: Five datasets focusing on dialects from Algeria, Egypt, Morocco, and the Gulf regions Nagoudi et al. (2023b). Classification: Five datasets for binary, regional, and country-level dialect identification Abdul-Mageed et al. (2021d, 2024); Elmadany et al. (2022); Abdul-Mageed et al. (2021b); Ahmed et al. (2024). STS: A novel synthetic dataset for Egyptian text similarity generated using Command-R+.
@@ 4.3 Domain-specific ArabicMTEB
Arabic Text retrieval tasks are currently trending in real-world applications. They are utilized across multiple fields, including healthcare, finance, and legal sectors. Having specialized evaluation datasets is crucial for building text embeddings tailored to these domains. To address this need, we propose Domain-specific ArabicMTEB, a fork of the larger ArabicMTEB benchmark. Domain-specific ArabicMTEB specializes in news, finance, legal, medical, and general knowledge domains, representing a close approximation to real-world scenarios. The creation of this benchmark involved collecting Arabic documents from these specialized sources and from Arabic Wikipedia. We split and chunked these documents into texts of 1,024 tokens in length. Subsequently, we randomly selected chunks and employed GPT4-Turbo OpenAI et al. (2024) to generate five different styles of queries for each chunk. We filtered out duplicate and repeated queries using GPT3.5 OpenAI et al. (2024) to ensure a high-quality evaluation dataset. Our evaluation data creation pipeline is visualized in Figure 2. The resulting benchmark, which we call ArabicMTEB Lite, contains $10k$ queries and $100k$ documents spanning the domains described above.
@@ 4.4 Cultural ArabicMTEB
To show that our models are culturally aware, we have introduced Cultural ArabicMTEB, a collection of datasets from 20 different Arab countries where we focus on specific cultural aspects like their geography, history, etc. To construct the Cultural ArabicMTEB, we use Arabic Wikipedia as our primary data source. For each included Arab country, we extract articles related to that country from its corresponding Wikipedia portal. The portal covers multiple categories (e.g., geography, economy, history) with subcategories (e.g., local movies, food items). This process resulted in $5K$ to $55K$ articles per country. The next step was to generate retrieval questions and passages for each country. For this, we use GPT-4o-mini to develop, for each passage (from an article), a corresponding question whose specific answer is available within the passage itself. Following the same methodology but applied to Egyptian and Moroccan dialectal versions of Wikipedia, we generate dialectal queries and their corresponding passages using Command-R+. Cultural ArabicMTEB  contains $1k$ queries and an average of $15k$ documents from various countries as described above.
@@ 5 Evaluation
Table 4: Overall ArabicMTEB results
@@
Model News Legal Medical Finance Wikipedia Avg Cost $\downarrow$
@@
Model MSA-Culture Egyptian-DIA Morocco-DIA Avg.
@@
Table 8: The impact of Synthetic Data on Swan performance. ArRTR: Arabic retrieval, DOM-RTR: Domain-specific retrieval, and DIA-RTR: Dialectal Retrieval
We evaluate the performance of our models, Swan-Small and Swan-Large, across the multiple proposed ArabicMTEB benchmarks and compare them with existing state-of-the-art models, including MARBERT Abdul-Mageed et al. (2020a), ARBERTv2 Elmadany et al. (2022), CamelBERT Inoue et al. (2021), multilingual E5 models Wang et al. (2024b), and Arabic-triplet-Matryoshka-V2 (ATM-V2) Nacar and Koubaa (2024). The evaluation results encompasses overall ArabicMTEB  (Table 4), dialectal ArabicMTEB (Table 5), domain-specific ArabicMTEB tasks (Table 6), and cultural ArabicMTEB (Table 7). In these tables, the tasks will be referred to as RTR: Retrieval, STS: Semantic Textual Similarity, PairCLF: Pair Classification, CLF: Classification, CLR: Clustering, RRK: Reranking, and BTM: BiText Mining.
@@
Synthetic Data Analysis We systematically analyze the impact of synthetic data on the performance of Swan-Small and Swan-Large using different combinations of training datasets. Table 8 presents the results for the base models, models trained with additional human-generated Arabic data, and models enhanced using synthetic subsets such as MSA, domain-specific, and dialectal data. When comparing the initial Swan-Small (average score of 32.46) to its version trained with synthetic MSA data, we observe a significant increase in average performance to 48.42, representing an improvement of more than 16 points. Similarly, Swan-Large benefits from a 6.52-point boost in average performance (from 55.39 to 61.91) with the inclusion of synthetic MSA data. Inclusion of Synthetic data in Although synthetic data significantly enhances model performance, it can introduce biases due to the reliance on specific patterns in the generated content. We ensured our synthetic data generation diversity by varying the data sources and generating dialectal data for multiple regions, including Egypt, Morocco, and the Gulf states, to mitigate this. We also analysed our models by examining whether MSA data received higher accuracies in retrieval tasks in Table 8. Further, our synthetic data generation pipeline was subjected to human verification for correctness and balance across cultural contexts. Although synthetic data significantly enhances model performance, it can introduce biases due to the reliance on specific patterns in the generated content. We ensured our synthetic data generation diversity by varying the data sources and generating dialectal data for multiple regions, including Egypt, Morocco, and the Gulf states, to mitigate this. We also analysed our models by examining whether MSA data received higher accuracies in retrieval tasks in Table 8. Further, our synthetic data generation pipeline was subjected to human verification for correctness and balance across cultural contexts.
@@ 6 Conclusion
In this paper, we introduced Swan-Small and Swan-Large, along with the comprehensive ArabicMTEB  benchmark for evaluating Arabic text embeddings. Our models demonstrate outstanding performance, benefiting from the strategic use of hard negatives and synthetic data in training. The evaluation across multiple benchmarks demonstrates that both Swan-Small and Swan-Large set new standards in Arabic-centric NLP tasks. They outperform existing state-of-the-art models in both cross-lingual and Arabic-specific tasks while being cost-effective and capable of understanding cultural context—making them ideal for real-world applications in diverse Arabic language settings.
@@ 7 Limitations
While the development of the Swan  models and the introduction of the ArabicMTEB benchmark mark significant advancements in Arabic text embeddings, there are some limitations to consider.
The societal implications of deploying dialect-aware models, such as Swan, require careful consideration. While these models can bridge gaps in NLP for Arabic-speaking regions, there is a risk of inadvertently reinforcing biases or language hierarchies, particularly in areas where particular dialects are stigmatized or underrepresented. For instance, users in communities with dialects associated with lower socioeconomic status may feel marginalized if their dialect is not adequately supported. To mitigate these concerns, we have prioritized the inclusion of low-resource dialects and ensured that our synthetic data generation pipeline accounts for dialectal diversity. Additionally, future versions of our models will include further dialectal balancing, specifically focusing on underrepresented communities.
@@ 8 Ethical Statement
All research and development activities for the Swan models and ArabicMTEB benchmark were conducted with a commitment to ethical standards. Data collection and usage adhered to privacy and confidentiality norms, ensuring no sensitive information was utilized without proper anonymization and consent. We acknowledge the potential biases introduced by synthetic data and have taken steps to mitigate these through diverse data sources and rigorous evaluation.
@@ Appendix A ArMistral Training
Pretraining data We further pretrain it on a large and diverse Arabic dataset, including all categories of Arabic, namely Classical Arabic (CA), Dialectal Arabic (DA), and MSA. This data is aggregated from various sources: AraNewsv2 Nagoudi et al. (2020), El-Khair El-Khair (2016), Gigaword, [^2] LDC Catalog Link OSCAR Suárez et al. (2019), OSIAN Zeroual et al. (2019), 101 Billion arabic words Aloui et al. (2024a), Wikipedia Arabic, and Hindawi Books. [^3] OpenITI corpus (v1.6). We also derived ArabicWeb22 (A) and (B) from the open source Arabic text 2022.444ArabicText-2022 data This pretraining dataset was cleaned, filtered and deduplicated using Bhatia (2023). We have also ensured that the model is pretrained in multiple domains, enhancing its results as seen in Table 9.
@@
Alignment Dataset We collected an alignment dataset from Quora and Mawdoo websites and then we took the gold answers as the choosen and we generated the rejected using AceGPT-7B Huang et al. (2024).
Results
@@
Table 9: Comparison of ArMistral with other Arabic LLMs
As seen from Table 9, Our ArMistral-Chat model outperforms all existing Arabic LLMs.
Appendix B Training methodology
Here, “{task_instruction}” refers to a one-sentence description of the embedding task taken from Table 11, which outlines the instructions for different tasks. Using a pretrained large language model (LLM), we append a [EOS] token at the end of both the modified query and the document. These are then input into the LLM to extract embeddings $\mathbf{h}_{q^{+}_{\text{inst}}}$ and $\mathbf{h}_{d^{+}}$ from the vector at the last [EOS] layer. The training of the embedding model is conducted using the InfoNCE loss function van den Oord et al. (2019b), which is widely recognized for its effectiveness in learning high-quality embeddings. The objective is minimized using the following formulation:
Appendix C Datasets overview
@@
Appendix D Polydedupe: versatile cleaning Pipeline
@@
Table 10: Benchmark Datasets Overview. Abbreviations: S2S = Sentence to Sentence, S2P = Sentence to Paragraph, P2P = Paragraph to Paragraph.
Appendix E Prompts for evaluation
@@
Appendix F Full Leaderboard
Table 12 presents the performance comparison of various models on different tasks within the ArMTEB benchmark. It includes metrics for Retrieval, Semantic Textual Similarity (STS), Pair Classification (PairCLF), Classification (CLF), Re-ranking, Clustering, and Bitext Mining (BTM). The table lists each model, its dimensionality, and the scores for each task, along with an overall average score. The results highlight the strengths and weaknesses of each model across a range of tasks, providing a comprehensive overview of their performance.
@@ Appendix F Full Leaderboard
multilingual-e5-large 1024 64.01 59.45 75.06 53.43 70.79 42.49 66.33 61.65
e5-mistral-7b-instruct 4096 56.34 57.02 70.24 53.21 66.24 39.44 70.50 59.00
Swan-Base 768 58.42 58.44 74.93 57.34 68.43 40.43 42.45 57.21
multilingual-e5-base 768 56.91 57.99 74.30 52.30 69.07 42.56 33.90 55.29
multilingual-e5-small 384 55.14 56.73 73.97 50.85 67.92 42.37 38.47 55.06
@@
Table 12: ArMTEB Results.
Appendix G Inference Latency.
@@
Inference latency is very critical in deploying machine learning models, especially in real-time applications with crucial response time. It refers to the time taken by a model to predict received input. In the context of text embedding models such as Swan-Base and Swan-Large, lower latency is particularly valuable for user-facing services that rely on fast processing of natural language input, such as chatbots and search engines. From Figure 4, we find that Swan-Large, despite its larger size indicated by a larger bubble, has optimized inference times due to architectural efficiencies, and Swan-Base strikes the perfect balance between size, performance, and latency. We compare the performance of the models from Table 4.
Appendix H Impact of Hard Negatives
Model (HN) 1 3 7 15 31
Swan-Small 48.84 52.19 54.13 56.25 51.93
Swan-Large 59.48 59.35 60.42 59.44 59.83
Table 13: Impact of number of Hard Negatives (HN).
Hard negatives are challenging examples that are nearly correct but ultimately incorrect, forcing the model to learn more nuanced distinctions between the different classes. The process involves converting all documents into a vector form within the embedding space. Subsequently, these document embeddings are compared using the cosine similarity score to establish their relevance to the query. Once all documents are scored, they are sorted by their similarity to the query. The top-ranked document is typically the positive example, while the rest are potential negatives. Our experiments assess the impact of varying the hard negatives used while training our models, Swan-Large and Swan-Base. We train each model with different quantities of hard negatives. Namely, we experiment with using 1, 3, 7, 15, and 31 hard negatives per training instance. Swan-Base reaches its highest performance at 56.25 with 15 hard negatives. This model shows a general upward trend in performance as the number of hard negatives increases, peaking at 15 but then declining slightly when the number increased to 31. This pattern suggests that while additional hard negatives initially provide beneficial learning challenges, there can be a point of diminishing returns where too much complexity hinders further learning. Swan-Large shows a peak in performance with $60.42$ when trained with seven hard negatives, indicating an optimal level of challenge that enhances learning without overwhelming the model. Interestingly, further increases in hard negatives do not improve performance, suggesting a threshold beyond which additional complexity does not translate to better learning outcomes.
Appendix I STS Dataset Creation:

### v3

@@ 4 ArabicMTEB Benchmark
Task Datasets Langs Dialects Metric
@@
In this section, we present ArabicMTEB, a comprehensive Arabic-centric text embedding benchmark designed to evaluate text embeddings across a wide range of tasks and scenarios. ArabicMTEB addresses the limitations of existing benchmarks that either exclude Arabic or lack coverage of diverse Arabic language varieties, dialects, and cultural nuances. Our benchmark includes $94$ datasets spanning $8$ distinct tasks, as summarized in Table 4. Further details about the datasets used in  benchmark  can be found in Appendix B. ArabicMTEB was developed to provide comprehensive coverage of Arabic text embedding capabilities, ensuring the inclusion of MSA and other varieties. It offers diverse task types, such as retrieval, classification, and semantic similarity, to evaluate embeddings holistically across different scenarios by incorporating novel domain-specific, dialectal, and country-level culturally aware datasets, ArabicMTEB represents a more applicable and realistic assessment of Arabic text embeddings.
@@ 4.1 Task Categories
ArabicMTEB categorizes evaluation datasets into the following key task categories, with each type providing a unique perspective on the capabilities of text embeddings. The corresponding metadata for each task, covering the considered number of datasets, number of languages, number of dialects, and evaluation metric, is presented in Table 4.
@@
Semantic Textual Similarity (STS). STS measures the correlation between the embeddings of two sentences, assessing their semantic similarity. ArabicMTEB includes five STS datasets like STS17 and STS22 Cer et al. (2017), along with two synthetic datasets generated using GPT-4 (Details of the creation of these datasets can be found in Appendix G).
@@
Clustering. Clustering groups sentences into clusters based on embedding similarity, evaluating unsupervised learning performance. ArabicMTEB includes four clustering datasets, such as Arabic News Articles and stance detection datasets from  Baly et al. (2018).
@@ 4.2 Dialectal ArabicMTEB
Dialectal ArabicMTEB is a specialized fork of the original ArabicMTEB, focusing exclusively on Arabic dialectal datasets. This extension addresses the unique challenges posed by the significant variations in Arabic dialects across different regions, which have been underrepresented in NLP research. While research on dialectal Arabic text embedding has been limited, dialectal ArabicMTEB fills this gap by providing a comprehensive collection of $19$ datasets specifically curated to evaluate embeddings’ performance on diverse Arabic dialects. These datasets span multiple tasks, offering a robust framework for assessing model performance across various dialectal contexts: (1) Bitext Mining. Eight datasets covering dialects such as Algerian, Egyptian, Jordanian, Lebanese, Moroccan, Saudi, and Yemeni Nagoudi et al. (2022); Bouamor et al. (2014). (2) Retrieval. Five datasets focusing on dialects from Algeria, Egypt, Morocco, and the Gulf regions Nagoudi et al. (2023b). (3) Classification: Five datasets for binary, regional, and country-level dialect identification Abdul-Mageed et al. (2021d, 2024b); Elmadany et al. (2022); Abdul-Mageed et al. (2021b); Ahmed et al. (2024). (4) STS. A novel synthetic dataset for Egyptian text similarity generated using Command-R+.
@@ 4.3 Domain-Specific ArabicMTEB
Arabic Text retrieval tasks are currently trending in real-world applications. They are utilized across multiple fields, including healthcare, finance, and legal sectors. Having specialized evaluation datasets is crucial for building text embeddings tailored to these domains. To meet this need, we introduce domain-specific ArabicMTEB, a specialized fork of the broader ArabicMTEB benchmark. Domain-specific ArabicMTEB focuses on the news, finance, legal, medical, and general knowledge domains, offering a closer approximation to real-world scenarios. The creation of this benchmark involves collecting Arabic documents from these specialized sources and from Arabic Wikipedia. We then segment and chunk the documents into texts of $1,024$ tokens each. Subsequently, we randomly select chunks and employ GPT4-Turbo OpenAI et al. (2024) to generate five different styles of queries for each chunk. We filter out duplicate and repeated queries using GPT3.5 OpenAI et al. (2024) to ensure a high-quality evaluation dataset. Our evaluation data creation pipeline is visualized in Figure 2. The resulting benchmark, which we call ArabicMTEB-Lite, contains $10k$ queries and $100k$ documents spanning the domains described above.
@@ 4.4 Cultural ArabicMTEB
To show that our models are culturally aware, we have introduced Cultural ArabicMTEB, a collection of datasets from 20 different Arab countries where we focus on specific cultural aspects like their geography, history, etc. To construct the Cultural ArabicMTEB, we use Arabic Wikipedia as our primary data source. For each included Arab country, we extract articles related to that country from its corresponding Wikipedia portal. The portal covers multiple categories (e.g., geography, economy, history) with subcategories (e.g., local movies, food items). This process resulted in $5$K to $55$K articles per country. Next, we generate retrieval questions and passages for each country. For this, we use GPT-4o-mini to develop, for each passage (from an article), a corresponding question whose specific answer is available within the passage itself. Following the same methodology, but applied to Egyptian and Moroccan dialectal versions of Wikipedia, we generate dialectal queries and their corresponding passages using Command-R+. Cultural ArabicMTEB  contains $1k$ queries and an average of $15k$ documents from various countries as described above.
@@ 5 Evaluation
Table 5: Overall ArabicMTEB results.
@@
Model News Legal Med Fin Wiki Avg Cost $\downarrow$
@@
Model MSA-Cult Egy-DIA Mor-DIA Avg.
@@
Table 9: The impact of Synthetic Data on Swan performance. ArRTR: Arabic retrieval, DOM-RTR: Domain-specific retrieval, and DIA-RTR: Dialectal Retrieval.
We evaluate the performance of our models, Swan-Small and Swan-Large, across the multiple proposed ArabicMTEB benchmarks and compare them with existing SoTA models, including MARBERT Abdul-Mageed et al. (2020a), ARBERTv2 Elmadany et al. (2022), CamelBERT Inoue et al. (2021), multilingual E5 models Wang et al. (2024b), and Arabic-triplet-Matryoshka-V2 (ATM-V2) Nacar and Koubaa (2024). Our evaluation results encompasses overall ArabicMTEB  (Table 5), dialectal ArabicMTEB (Table 6), domain-specific ArabicMTEB tasks (Table 7), and cultural ArabicMTEB (Table 8). In these tables, the tasks will be referred to as RTR: Retrieval, STS: Semantic Textual Similarity, PairCLF: Pair Classification, CLF: Classification, CLR: Clustering, RRK: Reranking, and BTM: BiText Mining.
@@
Synthetic Data Analysis. We systematically analyze the impact of synthetic data on the performance of Swan-Small and Swan-Large using different combinations of training datasets. Table 9 presents the results for the base models, models trained with additional human-generated Arabic data, and models enhanced using synthetic subsets such as MSA, domain-specific, and dialectal data. When comparing the initial Swan-Small (average score of 32.46) to its version trained with synthetic MSA data, we observe a significant increase in average performance to 48.42, representing an improvement of more than 16 points. Similarly, Swan-Large benefits from a 6.52-point boost in average performance (from 55.39 to 61.91) with the inclusion of synthetic MSA data.
@@ 6 Conclusion
In this paper, we introduced Swan-Small and Swan-Large, along with the comprehensive ArabicMTEB  benchmark for evaluating Arabic text embeddings. Our models demonstrate outstanding performance, benefiting from the strategic use of hard negatives and synthetic data in training. The evaluation across multiple benchmarks demonstrates that both Swan-Small and Swan-Large set new standards in Arabic-centric NLP tasks. They outperform existing SoTA models in both cross-lingual and Arabic-specific tasks while being cost-effective and capable of understanding cultural context—making them ideal for real-world applications in diverse Arabic language settings.
@@ 7 Limitations
While the development of the Swan models and the introduction of ArabicMTEB mark significant advancements in Arabic text embeddings, there are a number of limitations to consider. For example, although synthetic data significantly enhances model performance, it can introduce biases due to the reliance on specific patterns in the generated content. We ensured our synthetic data generation diversity by varying the data sources and generating dialectal data for multiple regions, including Egypt, Morocco, and the Gulf states, to mitigate this. We also analyzed our models by examining whether MSA data received higher accuracies in retrieval tasks in Table 9. Further, our synthetic data generation pipeline was subjected to human verification for correctness and balance across cultural contexts.
@@ 8 Ethical Statement
The societal implications of deploying dialect-aware models, such as Swan, require careful consideration. While these models can bridge gaps in NLP for Arabic-speaking regions, there is a risk of inadvertently reinforcing biases or language hierarchies, particularly in areas where particular dialects are stigmatized or underrepresented. For instance, users in communities with dialects associated with lower socioeconomic status may feel marginalized if their dialect is not adequately supported. To mitigate these concerns, we have prioritized the inclusion of low-resource dialects and ensured that our synthetic data generation pipeline accounts for dialectal diversity. Additionally, future versions of models should include further dialectal balancing, specifically focusing on underrepresented communities.
Importantly, all research and development activities for the Swan models and ArabicMTEB benchmark were conducted with a commitment to ethical standards. Data collection and usage adhered to privacy and confidentiality norms, ensuring no sensitive information was utilized without proper anonymization and consent.
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^4] https://alliancecan.ca and UBC Advanced Research Computing-Sockeye. [^5] https://arc.ubc.ca/ubc-arc-sockeye
@@ Appendix A ArMistral Training
Pretraining data We further pretrain it on a large and diverse Arabic dataset, including all categories of Arabic, namely Classical Arabic (CA), Dialectal Arabic (DA), and MSA. This data is aggregated from various sources: AraNewsv2 Nagoudi et al. (2020), El-Khair El-Khair (2016), Gigaword, [^6] LDC Catalog Link OSCAR Suárez et al. (2019), OSIAN Zeroual et al. (2019), 101 Billion arabic words Aloui et al. (2024a), Wikipedia Arabic, and Hindawi Books. [^7] OpenITI corpus (v1.6). We also derived ArabicWeb22 (A) and (B) from the open source Arabic text 2022.888ArabicText-2022 data This pretraining dataset was cleaned, filtered and deduplicated using Bhatia (2023). We have also ensured that the model is pretrained in multiple domains, enhancing its results as seen in Table 10.
@@
Alignment Dataset. We collected an alignment dataset from Quora and Mawdoo websites and then we took the gold answers as the choosen and we generated the rejected using AceGPT-7B Huang et al. (2024).
Results. As seen from Table 10, Our ArMistral-Chat model outperforms all existing Arabic LLMs.
@@
Table 10: Comparison of ArMistral with other Arabic LLMs.
Appendix B Datasets overview
@@
Appendix C Polydedupe: versatile cleaning Pipeline
@@
Table 11: Overview of ArabicMTEB datasets. S2S: Sentence to Sentence. S2P: Sentence to Paragraph. P2P: Paragraph to Paragraph.
Appendix D Prompts for evaluation
@@
Appendix E Full Leaderboard
Table 13 presents the performance comparison of various models on different tasks within the ArabicMTEB benchmark. It includes metrics for Retrieval, Semantic Textual Similarity (STS), Pair Classification (PairCLF), Classification (CLF), Re-ranking, Clustering, and Bitext Mining (BTM). The table lists each model, its dimensionality, and the scores for each task, along with an overall average score. The results highlight the strengths and weaknesses of each model across a range of tasks, providing a comprehensive overview of their performance.
@@ Appendix E Full Leaderboard
multilingual-e5-lg 1024 64.01 59.45 75.06 53.43 70.79 42.49 66.33 61.65
e5-mistral-7b-inst 4096 56.34 57.02 70.24 53.21 66.24 39.44 70.50 59.00
Swan-Small 768 58.42 58.44 74.93 57.34 68.43 40.43 42.45 57.21
multiling-e5-b 768 56.91 57.99 74.30 52.30 69.07 42.56 33.90 55.29
multiling-e5-s 384 55.14 56.73 73.97 50.85 67.92 42.37 38.47 55.06
@@
Table 13: ArabicMTEB  Results.
Appendix F Inference Latency.
@@
Inference latency is very critical in deploying machine learning models, especially in real-time applications with crucial response time. It refers to the time taken by a model to predict received input. In the context of text embedding models such as Swan-Small and Swan-Large, lower latency is particularly valuable for user-facing services that rely on fast processing of natural language input, such as chatbots and search engines. From Figure 4, we find that Swan-Large, despite its larger size indicated by a larger bubble, has optimized inference times due to architectural efficiencies, and Swan-Small strikes the perfect balance between size, performance, and latency. We compare the performance of the models from Table 5.
Appendix G STS Dataset Creation: